OpenStudy (anonymous):

Find the perimeter and area of a right triangle if one leg measures 9 cm and the other leg measures 40 cm. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.

OpenStudy (anonymous):

\[9+40+\sqrt{9^2+40^2}=90 \]or\[9+40+9\text{*Sec}\left[\text{ArcTan}\left[\frac{40}{9}\right]\right]=90 \]